OpenStudy (anonymous):

Solve this equation by completing the square. 3x2 - 4x = -2

OpenStudy (anonymous):

\[\huge 3x^2 - 4x = -2 \] dividing both sides by 3 we find: \[\huge x^2 - \frac{4}{3}x = \frac{-2}{3} \] \[\huge \rightarrow x^2 -2 \times \frac{2}{3}x = \frac{-2}{3} \] Now it is clear that we have to add the square of (2/3) both sides hence we find: \[\huge \rightarrow x^2 -2 \times \frac{2}{3}x +(\frac{2}{3})^2= \frac{-2}{3} +(\frac{2}{3})^2\] \[\huge \rightarrow x^2 -2 \times \frac{2}{3}x +(\frac{2}{3})^2= \frac{-2}{3} +\frac{4}{9}\] \[\huge \rightarrow [x-(\frac{2}{3})]^2= \frac{-6}{9} +\frac{4}{9}\] \[\huge \rightarrow [x-(\frac{2}{3})]^2= \frac{-2}{9} \] Now work it out further.

OpenStudy (anonymous):

3x^2 - 4x = -2 you can use delta